OpenStudy (anonymous):

What are the x- and y-intercepts of the graph of y = x^2 + 3x - 10? A)x-intercepts: (-2, 0) and (5, 0) y-intercept: (0, 10) B)x-intercepts: (-2, 0) and (-5, 0) y-intercept: (0, 10) C)x-intercepts: (-2, 0) and (5, 0) y-intercept: (0, -10) D)x-intercepts: (2, 0) and (-5, 0) y-intercept: (0, -10)

OpenStudy (haseeb96):

for x intercept put y=0 in the given equation and get the answer and for y intercept put the x=0 in the given equation and get the answer

OpenStudy (haseeb96):

for X -intercept put y=0 then equation becomes x^2+3x-10=0 x^2 -2x +5x -10=0 x(x-2) + 5(x-2)=0 (x+5)(x-2)=0 either x+5=0 or x-2 =0 x=-5 or x=2 this is yur x intercept

OpenStudy (haseeb96):

and for y intercept put x =0 then equation becomes y=-10 this is yur y intercept

OpenStudy (anonymous):

ok its D! thank you:)

OpenStudy (haseeb96):

most welcome my dear @kellyconnelly